Provide information for the medication: Acetaminophen ; Order is for _________________; Child’s weight is 18 kg...

  1. Provide information for the medication: Acetaminophen ; Order is for _________________; Child’s weight is 18 kg safe
  • discuss the drug class, side effects, patient and family education and nursing considerations.
    • calculate the safe dose your patient should receive______________________ what is the maximum dose? _________________ (remember this is a child)

  • concentration is ____________   SO how many ml. of the acetaminophen does the nurse administer draw up into the syringe: _____________
  • The dilution range for the medication: ______________ SO what is the dilution volume for the med? _______________
  • What is the infusion time is? _____________________
Answer #1

1For 18 kg the order for acetaminophen is 160mg\5ml.

2.Acetaminophen belongs to a class of drugs called Analgesics and Antipyretics.

Side effects are rash, nausea, headache. Other important sideffects include hypersensitivity reactions, serious skin reactions, kidney damage, Anaemia and thrombocytopenia.

Nursing intervention

  • Assess for desired therapeutic effects.
  • Periodic monitoring of CBC, coagulation studies, electrolytes, glucose, lipids, hepatic function test, renal function test.
  • Report any weight gain over 1kg /24 hr.
  • Assess the weight periodically.

Patient teaching

  • Advice the patient to take tablets whole or crushed.
  • Patient needs to read medication label and follow fosage guidelines.
  • Caution to not exceed recommended dossge.
  • Advice to talk to doctor prior to taking any OTC medicine because they may contain acetaminophen.
  • Teach patient to recognize signs of hepatotoxicity bleeding easy bruising and malaise results from overdose.

3.usual dose of acetaminophen is 10 to 15 mg/kg

10×18= 180mg

15×18=270mg

Safe dose range is from 180mg to 270 mg

Maximum dose=270mg.

4.Desired dose=160mg

Dose at hand=270mg.

Volume=5ml

Desired amount/Amount on hand× volume

160÷270×5ml=3.0ml